OpenStudy (anonymous):

simplify x squared -x-6/xsquared-2x-3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (mathteacher1729):

Factor the numerator and then factor the denominator. Cancel like factors, and you're done. :) QUESTION: can you factor something like \(x^2-x-6\) by itself?

OpenStudy (anonymous):

= (x-3)(x+2) / (x-3)(x+1) Cancelling x-3 = (x+2)/(x+1)

OpenStudy (anonymous):

(2x to 7th y squared/4xy to the 3rd)cube

OpenStudy (anonymous):

thanks a lot